# Practical privacy metrics for synthetic data

### v1

@@
Practical privacy metrics for synthetic data
Gillian M Raab
[…]
University of Edinburgh
gillian.raab@ed.ac.uk
[…]
School of GeoSciences
University of Edinburgh
Drummond Street
Edinburgh EH8 9XP, Scotland, United Kingdom
E-mail:
URL: https://www.synthpop.org.uk/
@@ 1 Introduction
In his recent review of synthetic data (SD) methodology Reiter (2023) comments:
@@
”While there is need to examine disclosure risks in synthetic data, there is no standard for doing so, especially in fully SD. Instead, disclosure checks tend to be ad hoc"
@@
This is in contrast to the variety of measures of utility available for SD; see Raab et al. (2021) for a list of these. We hope that these new functions that have been addded to the synthpop package for R 111see www.synthpop.org.uk  [^2] The new version 1.8-1 is currently available on Github at https://github.com/gillian-raab/synthpop will allow those planning to release SD to explore its disclosure risk and, where necessary, reduce it. While utility measures must be chosen that are relevant to the intended use of the SD, disclosure measures must focus on the possible harm to the privacy of an individual or other unit whose data contributed to the creation of the SD. The evaluation of the disclosure risk from SD must relate to the context of its release (see Elliot et al. (2020) for a discussion of this). We cannot expect a fixed rule, for example that a criterion for release requires a value of some disclosure measure beyond a threshold. Instead, we expect those releasing data to use the disclosure measures to evaluate potential harm to the data subjects, or to the data custodians, from information in the released data.
@@
We expect that disclosure risks from SD to be lower than those from the original data; we will refer to the original data used for the synthesis as the Ground Truth (GT). But in some cases, e.g. a large data set with only a few categorical variables, the disclosure risk from the GT may already be low. To evaluate the protection from disclosure afforded by a synthesis method, the risks for SD must be compared with equivalent risks for the GT. We consider two types of disclosure risk:
@@
  • identity disclosure: This refers to the ability to identify individuals in the data from a set of known characteristics that we will refer to as keys. Identity disclosure may be less relevant for completely synthesized [^3] Complete or full synthesis is when all values of all variables are replaced by synthetic values. This is in contrast to incomplete or partial synthesis where only some variables are replaced. data because there is no one-to-one correspondence between records in the GT and SD. But it may still be of interest since it is an important factor for attribute disclosure.
@@
  • attribute disclosure:This refers to the ability to find out from the keys something, not previously known, for an attribute associated with a record in the GT data.
The measures we describe here are appropriate for fully SD where all items in all records are replaced by synthetic values. Different measures have been developed for partially SD Drechsler and Reiter (2009); Reiter and Mitra (2009). Note that our disclosure methods treat numeric variables, by default, as if they were categories unless ngroups_keys and/or ngroups_target(s) are set. In our first example the income variable has been grouped into 20 categories, but the other numerical variable (depress) with only 21 categories has been left as it is. The number of groups formed may differ from the parameter setting e.g. if there are fewer than numgroups distinct values. [^4] The code uses grouping options from functions in the classInt package..
@@
The disclosure risk posed by SD can be reduced by using techniques from statistical disclosure control (sdc), such as aggregation of categories, smoothing of numeric values or removal of replicated uniques. These methods can be used to reduce disclosure risk by modifying the GT data before synthesis, or the SD before its is released. Some such methods are already available in the synthpop package. These include categorising, top/bottom coding and smoothing for continuous variables, and the merging of small categories for factors. The removal of replicated uniques is another option available in synthpop.
@@
There have been many recent proposals for making SD sets comply with Differential Privacy (DP) Dwork et al. (2006). DP is a very strong privacy guarantee that protects against an intruder with arbitrary external information about the subjects in the data, except for the one whose privacy is being protected. This is an unrealistic assumption and DP SD has been shown to have unacceptably low utility in many cases Bowen and Liu (2020); Stadler et al. (2022). We will not discuss these methods here, but note that we could use the metrics proposed here to evaluate disclosure risks for DP SD.
@@ 2 A simple example
Here we illustrate the basic use of disclosure.summary. If the parameter targets is not specified, all the variables in the SD that are not part of keys are used as targets. The identity disclosure measures are UiO for GT and repU for SD, and for attribute disclosure Dorig for original and DiSCO for synthetic. These and other measures will be explained in Sections 3.2 and 3.3.
First, a subset of 9 variables are selected fron the SD2011 data (a survey on quality of life in Poland) that is available as part of the synthpop package. A single SD set is created by the default method in synthpop: cart for each conditional distribution. Note that the variable income has values -8 that indicate not applicable, and the synthesis allows for this. The SD object s1555an object of class synds has a component syn that is a single synthetic data set. The disclosure functions can also be used with SD created by other methods either as single synthetic data sets or lists of repeated syntheses from the same original. Here we select 4 keys that represent items that might be known for members of this sample, or of the Polish population in 2011. By default,the 5 other variables become the targets: depress(score for depression) income ls (life satisfaction) marital (marital status), workab (intention to work abroad). The second target, income, is grouped into 20 categories, plus the -8 category. We first use disclosure.summary to create a disclosure.summary object for these keys, and we print out its identity component.
@@
R> s1 <- syn(ods, seed = 8564, print.flag = FALSE, cont.na = list(income = -8))
R> t1 <- disclosure.summary(s1, ods, print.flag = FALSE, plot = FALSE,
+  keys = c("sex", "age", "region", "placesize"),   ngroups_targets = c(0,20,0,0,0))
@@
The measure $UiO$ (Unique in Original) shows that 48% of the GT records would have unique combinations of these 4 keys. The term, ”singling out" is used in data protection regulation for this type of attribute disclosure [^6] See for example its use in the UK Information Commissioner’s guidance on anonymisation here https://ico.org.uk/media/about-the-ico/documents/4018606/chapter-2-anonymisation-draft.pdf. For the SD $RepU$ tells us that almost 15% of the GT records would be unique in the GT and also unique in the SD. For attribute disclosure we can examine the results as either a table or a plot. Here we see in Figure 1 the plot that would have been generated if plot had been set to TRUE in the code above, the attribute disclosure results would have been plotted as shown in Figure 1. Details of how identity and disclosure measures are calculated can be found in Section 3. The results in the Figure 1 are in descending order by the attribute disclosure risk in the SD.
Figure 1: Plot from disclosure.summary with plot = TRUE, the default value.
The measure $Dorig$ tells us that these 4 keys would identify a unique value of each of these targets for over 50% of the GT records. We get lower levels for $DiSCO$, the proportion of the GT records that are disclosive in the GT and also in the SD with a correct attribution to the target. The first two variables shown in Figure 1 have additional labels that flag possible contributions to disclosure from knowledge of 1-way or 2-way relationships in the GT data. This is detailed in Section 4.
Previous work on attribute disclosure Elliot (2014); Taub et al. (2018) has used the Correct Attribution Probability ($DCAP$) as a disclosure measure for SD. This is calculated as the percentage of records correctly predicted from the synthetic data by a set of keys. However it has been suggested Chen et al. (2019), that it may be as much a measure of utility as disclosure risk and modifications have been suggested (Taub and Elliot (2019); Little et al. (2022)) one of which ($TCAP$) is close to our proposed measure $DiSCO$. For completeness the measures $DCAP$ and $TCAP$ are calculated by disclosure; see Appendix 2.
@@ 3.1 Setting the scene
These disclosure measures are intended to assess what a person who only has access to the SD can infer about known individuals who are present in the GT data. We will use the term "intruder" for such a person, though no malicious intent is implied. The intruder is assumed to have information for one or more individuals about the value of certain key variables that are present in the same format in the GT and the SD. They attempt first to see if the individual is present, and then to determine the value of other items in the data file that we refer to as targets. We are assuming a worst-case scenario where the intruder believes they are querying the GT data. [^7] This may not be too unrealistic if the data are made available inadvertently, or if the intruder thinks that efforts to label the SD as e.g. ”Fake Data” are thought just to be a cover up. It may also be a reqasonable measure to use without this scenario, since it compares the disclosiveness of the SD to that of th GT. Disclosure measures from the SD are each compared to similar measures for someone with access to the GT data. Here we will introduce the measures by an example. Formal definitions with notation and formulae are in Appendix 1. The first step in evaluating disclosure risk, as described here, is to identify a set of keys that might be expected to be known to an intruder. These keys are then combined to form a quasi-identifier that we designate as $q$. For example, if we have hospital records we might define age,sex date and hospital as keys and this would give a $q$ with levels such as ”78 | M | 1/1/2024 | WG" for a 78 year old man admitted to hospital WG on 1/1/2024.
@@ 3.2 Identity disclosure measures
The concept of k-anonymity is central to identity disclosure for microdata. First proposed in 1998 Samarati and Sweeney. (1998) it is discussed fully in Elliot et al. (2020). A table is k-anonymous if a record cannot be distinguished from at least k others. Thus 2-anonymous data will never identify just one individual. Based on this idea, the percentage of records for which the keys identify just one individual give identity disclosure measures. Tables of $q$ values are produced from the GT and the SD. $UiO$ and $UiS$ are the percentages of records with keys where the table value is 1. An intruder checking out a record for their known set of keys will look for it in the SD. Some records will not be in the SD and $UiOiS$ (Unique in GT in Synthetic) gives the % that would be found. These records are then checked for uniqueness in the synthetic data giving, $repU$ is the percentage of unique GT records that are also unique in the SD.
The percentage $repU$ has been used as a disclosure measure to evaluate SD by Jackson et al. (2022) and by Raab (2022)  [^8] Jackson et al. in Jackson et al. (2022) argue that the denominator for $repU$ should be $N_{s}$ rather than $N_{d}$. This is inappropriate because our scenario is to consider the risk to the GT data.. Replicated uniques are used in synthpop as part of the statistical disclosure control function, sdc, that includes the option of reducing disclosure risk by removing them from the SD. Nowok et al. Nowok et al. (2017) have evaluated this and give an example where this process has very little effect on utility. The function replicated.uniques [^9] For example by replicated.uniques (s2, ods, keys = c("sex","region","age","placesize") also calculates $repU$ using a different method from the one described here.
One of the outputs of the functions disclosure and disclosure.summary is ident, a table of identity disclosure measures as illustrated by this example, using the keys from the example in Section 2 but now calculated for a synthetic object with 5 data sets, using the one target depress. We first create t5 an object of class disclosure and then print out the identity and attribute disclosure measures for each synthetic data set.
The table for identity disclosure has $UiO$ and $repU$ as its first and last column. The 2nd and 3rd columns are $UiS$ calculated from the SD in the same manner as $UiO$ from the GT and $UiOiS$ the % of $UiO$ with $q$ that are in the SD but not necessarily unique. Each are steps towards calculating $repU$.
@@ 3.3 Attribute disclosure measures
We approach disclosure risk from the point of view of an intruder with access to the SD and to certain attributes (quasi-identifiers) known for one or more individuals in the GT data. The quasi-identifiers for the known record(s) can be combined to create a composite variable $q$ that can be created for all records in the GT and SD. Modelling what an intruder might do we calculate the following measures, each of whuch is a proportion of the GT records:
  • Look up $q$ for each GT record in the SD. The proportion found becomes $iS$ (in Synthetic).
@@
  • Then check if these apparent disclosures correspond to the value of $t$ in the GT data. The proportion of GT records for which this is true becomes $DiSCO$ Disclosive in Synthetic Correct Original.
Note that records contributing to $DiSCO$ may not be disclosive in the GT data, as this information would not be available to the intruder. A further measure $DiSDiO$ (Disclosive in Synthetic Disclosive in Original) restricts the score to those disclosive in the original. These measures are defined formally in Appendix 1.
@@
Here we print the table of attribute disclosure measures for the example in the previous section. As we move from $iS$ to $DiS$ and to $DiSCO$ we can see how the %disclosive is affected by different conditions. Here lack of $q$ levels in the SD retains just 64% of records. The requirement for a record to be disclosive in the SD reduces this to 34% and again to around 9% for those with a corect attribution. This reduces it to around 6% by restricting to records disclosive in the GT.
The $Dorig$ and $DiSCO$ measures are not restricted to disclosures that are identified from unique records for $q$ in either the GT or the SD. The number of records contributing to each disclosive $qt$ cell in the synthetic table is the denominator that applies to that record. The columns max denom and mean denom refer to the denominators for the records disclosive in the SD that contribute to $DiSCO$. We can see from the mean that here the majority of disclosive records had unique key combinations in the SD, and the maxima was 3 for the first synthesis and 4 for the others. Large denominators can be an indication that some of the disclosure may be coming from strong relationships between variables in the data that might even be expected a-priori. This aspect is discussed further in Section 4. The disclosure measures can be restricted to those with small denominators by using the parameter exclude_over_denom_lim to TRUE. The example in Section 2 is here run restricted to denominators of 1.
R> disclosure.summary(s1, ods, print.flag = FALSE, plot = FALSE,
@@
                     attrib.orig attrib.syn                   check1 Npairs
1 income                   48.38       3.36                               0
@@
3 ls                       48.38       9.84                               0
4 marital                  48.38      15.40                               0
5 workab 1way checks       48.38      20.48 Check  workab  level  NO      0
The $DiSCO$ values have decreased, as expected, but $UiO$ and $repU$ have increased and also that $Dorig$ is now the same for all targets and equal to the $UiO$ value before the denominator exclusion. This makes sense because removing large denominators from $UiO$ and $repU$ increases the number of uniques. Also, with denominators of 1, all unique values of $q$ are disclosive in the GT data.

### v2

@@
Abstract
This paper explains how synthpop package111see www.synthpop.org.uk  [^2] The new version 1.9-0 is currently available on Github at https://github.com/gillian-raab/synthpop has been extended to include functions to calculate identity and attribute disclosure risk measures for the original records. using information obtained from the corresponding synthetic data. The basic function disclosure calculates identity disclosure for a set of quasi-identifiers (keys) and attribute disclosure for one variable specified as a target from the same set of keys. The second function multi.disclosure is a wrapper for the first and it presents summary results for a set of targets. This short paper explains the measures of disclosure risk and documents how they are calculated. We recommend two measures: $RepU$ (replicated uniques) for identity disclosure and $DiSCO$ (disclosive in synthetic and correct in the original) for attribute disclosure. Both are expressed a % of the original records and both can be compared to similar measures calculated from the original data. Experience with using the functions on real data found that some apparent disclosures could be identified as coming from relationships in the data that would be expected to be known to anyone familiar with its features. We flag cases when this seems to have occurred and provide means of excluding them.
@@ 1 Introduction
In his recent review of thirty years of synthetic data Reiter [17] comments:
@@
”While there is need to examine disclosure risks in synthetic data, there is no standard for doing so, especially in fully synthetic data. Instead, disclosure checks tend to be ad hoc”
@@
This is in contrast to the variety of measures of utility available for synthetic data; see [16] for a list of these. While utility measures must be chosen that are relevant to the intended use of the synthetic data, disclosure measures must focus on the possible harm to the privacy of an individual or other unit whose data contributed to the creation of the synthetic data. The evaluation of the disclosure risk from synthetic data must relate to the context of its release (see [5] for a discussion of this). We cannot expect a fixed rule, for example that a criterion for release requires a value of some disclosure measure beyond a threshold. Instead, we expect those releasing data to use the disclosure measures to evaluate potential harm to the data subjects or to the data custodians from information in the released data. We hope that these new functions will allow the disclosure risk of synthetic data sets to be explored and, where necessary, reduced.
@@
We expect that disclosure risks from synthetic data to be lower than those from the original data. But in some cases, e.g. a large data set with only a few categorical variables, the disclosure risk from the original may already be low. To evaluate the disclosure risk of a synthesis method, the risks for synthetic data must be compared with equivalent risks for the original. We consider two types of disclosure risk:
@@
  • identity disclosure: This refers to the ability to identify individuals in the data from a set of known characteristics that we will refer to as keys. Identity disclosure may be less relevant for completely synthesized [^3] Complete or full synthesis is when all values of all variables are replaced by synthetic values. This is in contrast to incomplete or partial synthesis where only some variables are replaced. data because there is no one-to-one correspondence between records in the original and synthetic data sets. But it may still be of interest since it is an important factor for attribute disclosure.
@@
  • attribute disclosure:This refers to the ability to find out from the keys something, not previously known, for an attribute associated with a record in the original data.
The measures we describe here are appropriate for fully synthetic data where all items in all records are replaced by synthetic values. Different measures have been developed for partially synthetic data [3, 18]
@@
The disclosure risk posed by synthetic data can be reduced by using techniques from statistical disclosure control (sdc), such as aggregation of categories, smoothing of numeric values or removal of replicated uniques. These methods can be used to reduce disclosure risk by modifying the original data before synthesis, or the synthetic data before its is released. Some such methods are already available in the synthpop package. These include categorising, top/bottom coding and smoothing for continuous variables, and the merging of small categories for factors. The removal of replicated uniques is another option available.
@@
There has been many recent proposals for making synthetic data sets Differential Private (DP) (selection of refs). DP is a very strong privacy guarantee that protects against an intruder with arbitrary external information about the subjects in the data, except for the one whose privacy is being protected. This is an unrealistic assumption and DP synthetic data has been shown to have unacceptably low utility in many cases [1, 20]. We will not discuss these methods here, but note that we could use the metrics proposed here to evaluate disclosure risks for DP synthetic data.
@@ 2 A simple example
Here we illustrate the basic use of multi.disclosure. If the parameter targets is not specified, all the variables in the synthetic data that are not part of keys are used as targets. The identity disclosure measures are UiO for original and repU for synthetic, and for attribute disclosure Dorig for original and DiSCO for synthetic. These and other measures will be explained in Sections 3.2 and 3.3.
First, a subset of 9 variables are selected fron the SD2011 data (a survey on quality of life in Poland) that is available as part of the synthpop package. A single synthetic data set is created by the default method in synthpop: cart for each conditional distribution. The synthetic data object s1444an object of class synds has a component syn that is a single synthetic data set. The disclosure functions can also be used with synthetic data created by other methods either as single synthetic data sets or lists of repeated syntheses from the same original. Here we select 4 keys that represent items that might be known for members of this sample, or of the Polish population in 2011. By default,the 5 other variables become the targets: income (rounded to give 406 distinct values) marital (marital status), workab (intention to work abroad), ls and depress (scores for life satisfaction and depression with 7 and 21 distinct values). We first use multi.disclosure to create a multi.disclosure object for these keys, and we print out its identity component.
@@
R> s1 <- syn(ods, seed = 8564, print.flag = FALSE)
R> t1 <- multi.disclosure(s1, ods, print.flag = FALSE, plot = FALSE,
+     keys = c("sex", "age", "region", "placesize"))
@@
The measure $UiO$ (Unique in Original) shows that 48% of the original records would have unique combinations of these 4 keys. The term, ”singling out” is used in data protection regulation for this type of attribute disclosure [^5] See for example its use in the UK Information Commissioner’s guidance on anonymisation here https://ico.org.uk/media/about-the-ico/documents/4018606/chapter-2-anonymisation-draft.pdf. For the synthetic data $RepU$ tells us that almost 15% of the original records would be unique in the original and also unique in the synthetic data.
For attribute disclosure we can examine the results as either a table or a plot. Here we see in Figure 1 the plot that would have been generated if plot had been set to TRUE in the code above, the attribute disclosure results would have been plotted as shown in Figure 1. Details of how identity and disclosure measures are calculated can be found in Section 3.
Figure 1: Plot from multi.disclosure with plot = TRUE, the default value.
The results in the plot are in descending order by the disclosure in the synthetic data. The measure $Dorig$ tells us that these 4 keys would identify a unique value of each of these targets for over 50% of the original records. $DiSCO$, the proportion of the original records that are disclosive in the original and also in the synthetic data with a correct attribution to the target give lower values. The first two variables shown in Figure 1 have additional labels that flag possible contributions to disclosure from knowledge of 1-way or 2-way relationships in the original data. This is detailed in Section 4.
Previous work on attribute disclosure [4, 22] has used the Correct Attribution Probability ($DCAP$) as a disclosure measure for synthetic data. This is calculated as the percentage of records correctly predicted from the synethetic data by a set of keys. However it has been suggested [7, 2], that it may be as much a measure of utility as disclosure risk and modifications have been suggested [21, 10] one of which ($TCAP$) is close to our proposed measure, $DiSCO$. For completeness the measures $DCAP$ and $TCAP$ are calculated by disclosure; see Appendix 2.
@@ 3.1 Setting the scene
These disclosure measures are intended to assess what a person who only has access to the synthetic data can infer about known individuals who are present in the original data. We will use the term ”intruder” for such a person, though no malicious intent is implied. The intruder is assumed to have information for one or more individuals about the value of certain key variables that are present in the same format in the original and the synthetic data. They attempt first to see if the individual is present, and then to determine the value of other items in the data file that we refer to as targets. We are assuming a worst-case scenario where the intruder believes they are querying the original data. [^6] This may not be too unrealistic if the data are made available inadvertently, or if the intruder thinks that efforts to label the synthetic data as e.g. ”Fake Data” are thought just to be a cover up. Disclosure measures from the synthetic data are each compared to similar measures for someone with access to the original data. Here we will introduce the measures by an example. Formal definitions with notation and formulae are in Appendix 1. The first step in evaluating disclosure risk, as described here, is to identify a set of keys that might be expected to be known to an intruder. These keys are then combined to form a quasi-identifier that we designate as $q$. For example, if we have hospital records we might define age,sex date and hospital as keys and this would give a $q$ with levels such as ”78 | M | 1/1/2024 | WG” for a 78 year old man admitted to hospital WG on 1/1/2024.
@@ 3.2 Identity disclosure measures
The concept of k-anonymity is central to identity disclosure for microdata. First proposed in 1998 [19] it is discussed fully in [5]. A table is k-anonymous if a set of keys identifies at most $k-1$ individuals. Thus 2-anonymous data will never identify just one individual. Based on this idea, the percentage of records for which the keys identify just one individual give identity disclosure measures. A table of $q$ values is produced from the synthetic and the original data. $UiO$ and $UiS$ are the percentages of records with keys where the table value is 1. An intruder checking out a record for their known set of keys will look for it in the synthetic data. Some records will not be in the synthetic data and $UiOiS$ (Unique in Original in Synthetic) gives the % that would be found. These records are then checked for uniqueness in the synethetic data giving, $repU$ is the percentage of unique original records that are also unique in the synthetic data.
The percentage $repU$ has been used as a disclosure measure to evaluate synthetic data by [9] and by [15]  [^7] Jackson et al. in [9] argue that the denominator for $repU$ should be $N_{s}$ rather than $N_{d}$. This is inappropriate because our scenario is to consider the risk to the original data.. Replicated uniques are used in synthpop as part of the statistical disclosure control function, sdc, that includes the option of reducing disclosure risk by removing them from the synthetic data. Nowok et al. [13] have evaluated this and give an example where this process has very little effect on utility. The function replicated.uniques [^8] For example by replicated.uniques (s2, ods, keys = c("sex","region","age","placesize") also calculates $repU$ using a different method from the one described here.
One of the outputs of the functions disclosure and multi.disclosure is ident, a table of identity disclosure measures as illustrated by this example, using the keys from the example in Section 2 but now calculated for a synthetic object with 5 data sets and using the one target depress. We first create t5 an object of class disclosure and then print out the identity and attribute disclosure measures for each synthetic data set.
The table for identity disclosure has $UiO$ and $repU$ as its first and last column. The 2nd and 3rd columns are $UiS$ calculated from the synthetic data in the same manner as $UiO$ from the original and $UiOiS$ the % of $UiO$ with $q$ that are in the synthetic data but not necessarily unique. Each are steps towards calculating $repU$.
@@ 3.3 Attribute disclosure measures
We approach disclosure risk from the point of view of an intruder with access to the synthetic data and to certain attributes (quasi-identifiers) known for one or more individuals in the original data. The quasi-identifiers for the known record(s) can be combined to create a composite variable $q$ that can be created for all records in the original and synthetic data. Modelling what an intruder might do we calculate the following measures, each of whuch is a proportion of the original records:
  • Look up $q$ for each original record in the synthetic data. The proportion found becomes $iS$ (in Synthetic).
@@
  • Then check if these apparent disclosures corresponds to the value of $t$ in the original data. The proportion of original records for which this is true becomes $DiSCO$ Disclosive in Synthetic Correct in Original.
Note that records contributing to $DiSCO$ may not be disclosive in the original data, as this information would not be available to the intruder. A further measure $DiSDiO$ (Disclosive in Synthetic Disclosive in Original) restricts the score to those disclosive in the original. These measures are defined formally in Appendix 1.
@@
Here we print the table of attribute disclosure measures for the example in the previous section. As we move from $iS$ to $DiS$ and to $DiSCO$ we can see how the %disclosive is affected by different conditions. Here lack of $q$ levels in the synthetic data retains just 64% of records. The requirement for a record to be disclosive in the synthetic data reduces this to 34% and again to around 9% for those with a corect attribution. This reduces it to around 6% by restricting to records disclosive in the original.
The $Dorig$ and $DiSCO$ measures are not restricted to disclosures that are identified from unique records for $q$ in either the original or the synthetic data. The number of records contributing to each disclosive $qt$ cell in the synthetic table is the denominator that applies to that record. The columns max denom and mean denom refer to the denominators for the records disclosive in the synthei(c data that contribute to the $DiSCO$ or $DISCDO$ measures. We can see from the mean that here the majority of disclosive records had unique key combinations in the synthetic data, and the maxima was 3 for the first synthesis and 4 for the others. Large denominators can be an indication that some of the disclosure may be coming from strong relationships between variables in the data that might even be expected a-priori. This aspect is discussed further in Section 4. The disclosure measures can be restricted to those with small denominators by using the parameter exclude_over_denom_lim to TRUE. The example in Section 2 is here run restricted to denominators of 1.
R> multi.disclosure(s1, ods, print.flag = FALSE, plot = FALSE,
@@
                     attrib.orig attrib.syn check1 Npairs check2
1 income                   48.38       4.02             0
@@
3 ls                       48.38       9.74             0
4 marital                  48.38      15.72             0
5 workab 1way checks       48.38      20.62     NO      0
The $DiSCO$ values have decreased, as expected, but $UiO$ and $repU$ have increased and also that $Dorig$ is now the same for all targets and equal to the $UiO$ value before the denominator exclusion. This makes sense because removing large denominators from $UiO$ and $repU$ increases the number of uniques. Also, with denominators of 1, all unique values of $q$ are disclosive in the original data.